# SynthFormer: Equivariant Pharmacophore-based Generation of Molecules for Ligand-Based Drug Design

### v1

@@ Abstract
Drug discovery is a complex and resource-intensive process, with significant time and cost investments required to bring new medicines to patients. Recent advancements in generative machine learning (ML) methods offer promising avenues to accelerate early-stage drug discovery by efficiently exploring chemical space. This paper addresses the gap between in silico generative approaches and practical in vitro methodologies, highlighting the need for their integration to optimize molecule discovery. We introduce SynthFormer, a novel ML model that utilizes a 3D equivariant encoder for pharmacophores to generate fully synthesizable molecules, constructed as synthetic trees. Unlike previous methods, SynthFormer incorporates 3D information and provides synthetic paths, enhancing its ability to produce molecules with good docking scores across various proteins. Our contributions include a new methodology for efficient chemical space exploration using 3D information, a novel architecture called Synthformer for translating 3D pharmacophore representations into molecules, and a meaningful embedding space that organizes reagents for drug discovery optimization. Synthformer generates molecules that dock well and enables effective late-stage optimization restricted by synthesis paths.
@@ 1 Introduction
Drug discovery is a complex, lengthy and expensive process (Wouters et al., 2020). Computer aided drug design has been shown to be effective (Schneider & Fechner, 2005), while generative machine learning (ML) methods offer a promising way to speed up the early stages of drug discovery by exploring molecules in an efficient and cost-effective way (Meyers et al., 2021). There are two main computational methods for early-stage drug discovery: de novo ligand design and virtual screening. State-of-the-art ML methods for molecule screening and generation, such as TargetDiff (Guan et al., 2023) and Pocket2mol (Peng et al., 2022), often overlook the practical aspects of synthesizing molecules in the lab. On the other hand, screening models like Equibind (Stärk et al., 2022) and Diffdock (Corso et al., 2023) are limited by the library that they need to screen. In vitro screening uses different methodologies. DNA encoded libraries (DEL) (Brenner & Lerner, 1992) and High-throughput screen (HTS) (Inglese et al., 2007) screening are costly, because they require the acquisition of large compound libraries, specialist equipment and use up large amounts of lab consumables. Dose-response assays (Crump et al., 1976) are limited by low throughput. Therefore, it is crucial to integrate in vitro technologies with in silico techniques to optimize molecule discovery, which will be specific for each target and disease to reduce cost and improve the speed of discovery.
For instance, DEL or dose-response screens can be enhanced with generative methods that suggest which molecules to screen, as each molecule would need to be ordered from a 3rd party supplier or made in house. These molecules are limited by the available building blocks and reactions (Fair et al., 2021). In contrast, HTS offers limited improvement potential from in silico methods, as its advantages stem from hardware optimization for a fixed setup and library. Consequently, not all combinations of ML and in vitro techniques provide acceleration; therefore, we should focus on identifying the right combination that can significantly enhance experimental outcomes.
@@
One key failure point for generative ML methods in drug discovery is synthesisability. Most methods rely on the Synthetic Accessibility (SA) score. However, it has been demonstrated that the SA score (Ertl & Schuffenhauer, 2009) is not a discriminative feature and is unable to determine whether a molecule is synthesizable, as shown in Synflownet (Cretu et al., 2024). There have been multiple works that use a generative tool for a bottom up approach (Gao et al., 2022) (Cretu et al., 2024). However, none of these models incorporate 3D information.
@@
Moreover, the most common molecule optimization approach relies on an external scorer blindly guiding the optimization as proposed by Guacamol (Brown et al., 2019) and Moses (Polykovskiy et al., 2018). However, in reality, the lead optimization process is heavily constrained by the available synthesis paths which also become the basis for patenting the Markush structure (Brown, 1991).
@@
  • We present a new methodology to encode the dataset to effectively learn how to explore the chemical space, based on 3D information.
  • We propose a novel architecture to translate from 3D pharmocophore representations to molecules, which we call Synthformer.
  • We demonstrate that Synthformer model is able to generate molecules that dock well.
  • We show that that the embedding space organises the reagents in a meaningful way allowing for a more drug discovery based optimization.
  • We propose a simple and a more effective optimization tool for later stage optimization which is restricted by synthesis paths.
@@ 2 Related work
Combinatorial Optimization. Approaches in combinatorial optimization employ a range of chemical building blocks and reaction templates to construct a combinatorial chemical space, such as Galileo (Meyenburg et al., 2023). These strategies then apply optimization techniques, such as genetic algorithms (Gao et al., 2021) and Monte Carlo tree search (Swanson et al., 2023), to explore this space for desired molecules. Prior to the advent of molecular deep learning, early efforts like SYNOPSIS (Vinkers et al., 2003) generated candidate molecules through virtual reactions, selecting them based on scoring functions. Recent advancements in deep learning for predicting chemical reactions (Coley et al., 2019) have further inspired methods that utilize neural networks to forecast reaction outcomes without relying on predefined templates.
@@
3D generative models. There are two primary approaches for shape-based generation: conditioning on a protein structure or on a known active ligand. Structure-based methods, such as TargetDiff (Guan et al., 2023) and Pocket2Mol (Peng et al., 2022), are designed to generate molecules that fit within a protein binding site. However, benchmark studies (Jocys et al., 2024) have demonstrated that while these methods can produce molecules that theoretically fit, they often fail to produce synthesis-ready compounds and may not bind as intended needing aditional modifications as shown in (Luo et al., 2024). Shape-based approaches, like SQUID (Adams & Coley, 2023) and LigDream (Skalic et al., 2019), that utilises a way to encode the shape or pharmacophores, and use another decoder to decode the molecules. Despite their promise, these networks have also struggled with generating synthesisable molecules.
Figure 1: We take the available protein structure or information, along with data from active ligands, which can typically be used to extract a hypothesis. However, in our case, we focus on the crystallized ligand, extracting its pharmacophores, and then use these pharmacophores as the basis to generate new compounds
@@
Molecule Optimization in a machine learning context was described in Guacamol (Brown et al., 2019) and Moses (Polykovskiy et al., 2018) as a process where molecules are improved by using scoring functions to evaluate how well they fulfill desired property profiles. The goal is to generate molecules that maximize these scores, while separating this optimization process from the challenge of selecting the appropriate scoring functions (Schneider & Fechner, 2005). This approach has been used in the ML community for SMILES based representation (Cao & Kipf, 2022), Graph based representation without 3D information, RL approaches (Cretu et al., 2024), as well as 3D approaches (Peng et al., 2022). At the same time, this optimization framing is detached from the drug discovery process in real life. In drug discovery, once HITs are identified and Structure-activity relationship is established, lead optimization involves exploiting the known synthetic pathways and exploring similar routes to improve the molecule. This was done in this lead optimization study (Harris & Faucher, 2019), where they explain how they went about finding and optimizing 4,5-Dihydropyrazoles as Mono-Kinase Selective, Orally Bioavailable and Efficacious Inhibitors of Receptor Interacting Protein 1 (RIP1) Kinase.
@@
Limitations of current state-of-the-art. Combinatorial optimization techniques are constrained by their reliance on predefined reaction templates, which result in a slow exploration of chemical space. On the other hand, deep learning models typically focus on only one part of the task: generating structurally and physically accurate structures or synthetically accessible compounds, but not both. Moreover, while deep learning methods excel in these areas, they often fail to integrate the optimization process with the specific needs of lead optimization. As a result, the generated compounds may not align well with the practical requirements of drug discovery, such as feasibility of synthesis or sufficient solubility for assay conditions. This detachment between structural design and practical application can hinder the overall success of deep learning approaches in optimizing drug candidates.
@@ 3 Method
3.1 Molecule Representation
Chemical space A chemical space $C$ is a set of molecules generated by a finite set of building block molecules $BB=\{B_{1},BB_{2},\ldots,BB_{N}\}\subset C$ and a set of reaction rules $R=\{R_{1},R_{2},\ldots,R_{M}\}$. Formally, a reaction rule maps reactants to a reaction product. For example, a reaction rule $R$ with two reactants can be denoted by:
$R:BB_{1}\times BB_{2}\rightarrow C$
$(BB_{1},BB_{2})\mapsto Y,$
where $BB_{1}$ and $BB_{2}$ are sets of molecules to which reaction $R$ can be applied, and $Y$ is the main reaction product. The chemical space is generated by starting from the building blocks and iteratively applying the reactions to every possible combination of molecules. There are some cases where a reaction has multiple possible main products. For ease of discussion, we assume that reactions produce only one main product in this section.
Every molecule in the chemical space is a product of a synthetic process, which involves applying reaction rules to building blocks and intermediate molecules to build a complex molecule. For example, $P2=R_{2}(P1,B_{3})$, where $P1=R_{1}(B1,B_{2})$ denotes a process in which we first apply reaction $R_{1}$ to building blocks $B_{1}$ and $B_{2}$, and then apply reaction $R_{2}$ to building block $B_{3}$ and the product of $R_{1}(B_{1},B_{2})$.
@@ 3.1 Molecule Representation
Pharmacophores are abstract representations of the molecular features necessary for molecular recognition of a ligand by a biological macromolecule, crucial for its biological activity. For each atom of the molecule that contains a pharmacophore we assign a one hot encoding. The one-hot encoding vector $\mathbf{p}=[p_{1},p_{2},p_{3},p_{4},p_{5},p_{6}]$ for the pharmacophore types available in the RDKit (Landrum, 2006) package represents the presence (1) or absence (0) of Hydrogen Bond Donors (HBD), Hydrogen Bond Acceptors (HBA), Aromatic Rings, Hydrophobic Centers, Positive Ionizable, and Negative Ionizable features, respectively.
3.2 Data preparation
3.2.1 Datasets
Reaction templates We use the publicly available reaction template set (Hartenfeller et al., 2012). The template set contains 58 reaction templates encoded in SMARTS string format.
Building blocks We use the building blocks from Mcule (Mcule Team, 2023). The building blocks that fail the RDKit sanitization check or do not match any reaction template are removed. Duplicate building blocks are also removed. The pre-processing procedure leads to the final building block set containing 10858 molecules.
3.2.2 Data generation
To generate the data, we begin by selecting a set of molecules and perform a sequence of $n$ random reactions from the available set, saving the resulting products. Next, we repeat the process but include the option to sample both the initial molecules and the previously generated products, effectively exploring the chemical space through random sampling. Once the molecules are sampled, we compute a random conformer for each one using Merck Molecular Force Field (MMFF) optimization from RDKit (Landrum, 2006). Afterwards, we extract the pharmacophores and 3D coordinates of each molecule. This results in an input dataset where the pharmacophores and 3D coordinates serve as the inputs, and the output prediction corresponds to the synthetic tree.
@@ 3.2.2 Data generation
Figure 2: The generation process begins by representing the pharmacophores as a fully connected graph, passed through a GNN to obtain pharmacophore embeddings. These embeddings are then propagated to a transformer decoder. The transformer decoder first receives a start token and predicts building block (BB1), followed by reaction 0. Next, it takes the fingerprint of BB1 as input to the decoder and predicts building block 2 (BB2), followed by reaction 1 (R1), which is applied to generate product 1 (P1). Product 1 is then used to predict building block 3 (BB3) and reaction 2 (R2), producing product 3(P3). This process repeats until the end token is generated.
Similarity scores To evaluate the similarity between the input molecule and the output molecule, we use the Tanimoto similarity score on three different fingerprints: (1) Morgan fingerprint of length 4096 and radius 2 (Morgan, 1965), (2) Morgan fingerprint of Murcko scaffold (Bemis & Murcko, 1996), and (3) Gobbi pharmacophore fingerprint (Gobbi & Poppinger, 1998). The three similarity scores indicate chemical similarities in three different aspects: overall structure, scaffold structure, and pharmacophore property, respectively, and are all normalized to [0, 1].

### v2

# SynthFormer: Equivariant Pharmacophore-based Generation of Synthesizable Molecules for Ligand-Based Drug Design

@@ Abstract
Drug discovery is a complex, resource-intensive process requiring significant time and cost to bring new medicines to patients. Many generative models aim to accelerate drug discovery, but few produce synthetically accessible molecules. Conversely, synthesis-focused models do not leverage the 3D information crucial for effective drug design. We introduce SynthFormer, a novel machine learning model that generates fully synthesizable molecules, structured as synthetic trees, by introducing both 3D information and pharmacophores as input. SynthFormer features a 3D equivariant graph neural network to encode pharmacophores, followed by a Transformer-based synthesis-aware decoding mechanism for constructing synthetic trees as a sequence of tokens. It is a first-of-its-kind approach that could provide capabilities for designing active molecules based on pharmacophores, exploring the local synthesizable chemical space around hit molecules and optimizing their properties. We demonstrate its effectiveness through various challenging tasks, including designing active compounds for a range of proteins, performing hit expansion and optimizing molecular properties.
Machine Learning, ICML
@@ 1 Introduction
Drug discovery is a complex, lengthy and expensive process (Wouters et al.,, 2020). Computer aided drug design has been shown to be effective (Schneider and Fechner,, 2005), while generative machine learning (GML) methods offer a promising way to speed up the early stages of drug discovery by exploring molecules in an efficient and cost-effective way (Meyers et al.,, 2021). There are two main computational methods for early-stage drug discovery: de novo ligand design and virtual screening. State-of-the-art ML methods for target based molecule generation, such as TargetDiff (Guan et al.,, 2023) and Pocket2mol (Peng et al.,, 2022), often overlook the practical aspects of synthesizing molecules in the lab. Conversely, screening models like Equibind (Stärk et al.,, 2022) and Diffdock (Corso et al.,, 2023) are limited by the libraries they must screen. In vitro screening is constrained by various factors, such as reactions, building blocks, and the methods used to obtain readings. DNA-encoded libraries (DELs) (Brenner and Lerner,, 1992), high-throughput screening (HTS) (Inglese et al.,, 2007), and Dose-Response Curves (Crump et al., 1976a, ) represent just a few of these approaches. HTS cannot be effectively optimized with machine learning, as its benefits arise from screening a fixed library, whereas DELs, while potentially optimized, are constrained by the building blocks and reactions used in their construction. Dose-response assays (Crump et al., 1976b, ) are limited by low throughput. Therefore, it is crucial to integrate in vitro technologies with in silico techniques to optimize molecule discovery for each target and disease, thereby reducing costs and improving the speed of discovery.
@@
One key failure point for generative ML methods in drug discovery is synthesisability. Most methods rely on the Synthetic Accessibility (SA) score. However, it has been demonstrated that the SA score (Ertl and Schuffenhauer,, 2009) is not a discriminative feature and is unable to determine whether a molecule is synthesizable, as shown in Synflownet (Cretu et al.,, 2024). There have been multiple works that use a generative tool for a bottom up approach (Gao et al.,, 2022) (Cretu et al.,, 2024). However, none of these models incorporate 3D information, which is essential for accurately understanding how molecules interact with their target protein, a critical factor in effective drug design.
@@
Moreover, the most common molecule optimization approach relies on an external scorer blindly guiding the optimization as proposed by Guacamol (Brown et al.,, 2019) and Moses (Polykovskiy et al.,, 2018). However, in reality, the lead optimization process is heavily constrained by the available synthesis paths which also become the basis for patenting the Markush structure (Brown,, 1991).
Pharmacophore-based ligand design has proven to be an effective approach for identifying early leads in drug discovery (Dash et al.,, 2019), particularly in cases where a protein structure with a defined binding mode is unavailable. When the structure of a protein target is not available or the mechanism of action is not well understood, protein folding algorithms, such as AlphaFold (Jumper,, 2021) or Boltz1 (Wohlwend et al.,, 2024), or homology modelling may not produce a model suitable for structure-based design. In these cases, scientists often rely on ligand-based design, although this approach is limited by the need for multiple ligands with affinity to the target protein as input. This underscores the need for models capable of generating diverse compound libraries that can be screened based on general hypotheses.
@@
  • We introduce the first methodology that incorporates both 3D information and synthesis to make pharmacophore ligand design possible.
  • We propose a novel Synthesis-Aware Decoder, called SynthFormer, for generating molecules by translating 3D pharmacophore representations into synthesizable compounds.
  • We demonstrate that the SynthFormer model is capable of generating molecules with strong docking performance, performing hit expansion, and can be used for molecule optimization.
@@ 2 Related work
Approaches in combinatorial optimization employ a range of chemical building blocks and reaction templates to construct a combinatorial chemical space, such as Galileo (Meyenburg et al.,, 2023). These strategies then apply optimization techniques, such as genetic algorithms (Gao et al.,, 2021) and Monte Carlo tree search (Swanson et al.,, 2023), to explore this space for desired molecules. Prior to the advent of molecular deep learning, early efforts like SYNOPSIS (Vinkers et al.,, 2003) generated candidate molecules through virtual reactions, selecting them based on scoring functions. Recent advancements in deep learning for predicting chemical reactions (Coley et al.,, 2019) have further inspired methods that utilize neural networks to forecast reaction outcomes without relying on predefined templates.
Synthesis-Aware Deep Learning
Related work spans various approaches to molecule design, based on reinforcement learning (RL) or self-supervised learning. RL-based methods, such as SynFlowNet (Cretu et al.,, 2024) and SynNet (Gao et al.,, 2022), optimize molecular properties by simulating sequential decisions, constrained by synthesizability and reaction feasibility. The core limitation is that the reward function has limited accuracy. The Synflownet model was guided by the docking module, which has an accuracy of 0.41. These methods rely on reward functions tailored to specific objectives, such as activity or docking scores. On the other hand, self-supervised learning approaches, like ChemProjector (Luo et al.,, 2024), learn meaningful molecular representations by predicting structural or functional transformations in a molecule-to-molecule framework. However, this does not allow to model activity cliffs (Zhang et al.,, 2023), as it is only transforming the molecules based on the graph and has no binding information.
3D Generative Models
@@
There are two primary approaches for shape-based generation: conditioning on a protein structure or on a known active ligand. Structure-based methods, such as TargetDiff (Guan et al.,, 2023) and Pocket2Mol (Peng et al.,, 2022), are designed to generate molecules that fit within a protein binding site. However, benchmark studies (Jocys et al.,, 2024) have demonstrated that while these methods can produce molecules that theoretically fit, they often fail to produce synthesis-ready compounds(Luo et al.,, 2024) and may not bind as intended. Ligand-based approaches, like SQUID (Adams and Coley,, 2023) and LigDream (Skalic et al.,, 2019), encode the shape or pharmacophores, and use another decoder to decode the molecules. Despite their promise, these networks have also struggled with generating synthesisable molecules.
Molecule Optimization
@@
Molecule optimization in a machine learning context, as described in Guacamol (Brown et al.,, 2019) and Moses (Polykovskiy et al.,, 2018), involves improving molecules by using scoring functions to evaluate how well they meet desired property profiles. The objective is to generate molecules that maximize these scores, while decoupling the optimization process from the challenge of selecting appropriate scoring functions (Schneider and Fechner,, 2005). This approach has been adopted by the ML community for various representations, including SMILES-based (Cao and Kipf,, 2022), graph-based without 3D information, reinforcement learning methods (Cretu et al.,, 2024), and 3D approaches (Peng et al.,, 2022). At the same time, this optimization framing is detached from the drug discovery process in real life. In drug discovery, once Hits are identified and Structure-Activity Relationship (SAR) is established, lead optimization involves exploiting the known synthetic pathways and exploring similar routes to improve the molecule. This was done in this lead optimization study (Harris and Faucher,, 2019), where they find and optimize inhibitors of RIP1 Kinase.
Limitations of Existing Methods
@@
Combinatorial optimization techniques are constrained by their reliance on predefined reaction templates, which result in a slow exploration of chemical space. On the other hand, deep learning models typically focus on only one part of the task: generating structurally and physically accurate structures or synthetically accessible compounds, but not both. Moreover, while deep learning methods excel in these areas, they often fail to integrate the optimization process with the specific needs of lead optimization. As a result, the generated compounds may not align well with the practical requirements of drug discovery, such as feasibility of synthesis or sufficient solubility for assay conditions. This detachment between structural design and practical application can hinder the overall success of deep learning approaches in optimizing drug candidates.
SynthFormer is a model designed to address these challenges. To the best of our knowledge this is the first model to translate 3D geometric profiles of pharmacophores and constructs molecules from available building blocks based on specified requirements, only operating in the plausible chemical space. Additionally, its architecture operates at the synthetic level, enabling an optimization process that mimics the work of a chemist in real drug discovery scenarios.
@@ 3 Method
SynthFormer introduces a novel methodology for molecule generation, characterized by innovative data preparation, novel input representations, and a unique architectural synergy between Equivariant Graph Neural Networks (EGNNs) and Synthesis-Aware Decoder (see Sec. 3.3 for both), as sketched in Figure 1. Our approach leverages a new type of input: desired pharmacophore features paired with their 3D coordinates. These inputs are encoded by the EGNN into rich embeddings that comprehensively capture critical spatial properties, providing a deeper understanding of molecular structure. These embeddings are subsequently processed by a transformer decoder, which autoregressively predicts a sequence of building blocks and reactions, ensuring a smooth and interpretable generation process.
To train and evaluate our model, we need to generate a custom dataset. Existing deep learning models for chemistry often operate at the atom level and rely on datasets like PDBBind (Wang et al.,, 2004), QM9 (Wu et al.,, 2018), or QMugs (Isert et al.,, 2022). However, since our approach focuses on pharmacophores and their synthetic trees—and because we rely on Enamine’s extensive chemical database (to which they graciously provided access)—it is necessary to construct our own dataset. This dataset will support the development and testing of the first model specifically designed to operate on pharmacophores in the context of synthetic tree generation, as elaborated in the following part.
3.1 Encoder Input: Pharmacophore
For data preparation, we construct input batches by combining pharmacophore features and 3D spatial information derived from conformers. Pharmacophores, represented as one-hot encoded vectors, are assigned to each atom, while the conformers provide the corresponding $(x,y,z)$ coordinates. Together, these are used to create the input data for the EGNN encoder, ensuring it captures both the spatial geometry and the pharmacophore profile of the molecule.
@@ 3.1 Encoder Input: Pharmacophore
Figure 1: The generation process begins by representing the pharmacophores as a fully connected graph, passed through an EGNN to obtain pharmacophore embeddings. These embeddings are then propagated to a transformer decoder. The transformer decoder first receives a start token and predicts building block $B_{1}$, followed by reaction $R_{0}$. Next, it takes the fingerprint of $B_{1}$ as input to the decoder and predicts building block $B_{2}$, followed by reaction $R_{1}$, which is applied to generate product $P_{1}$. $P_{1}$ is then used to predict building block $B_{3}$ and reaction $R_{2}$, producing product $P_{3}$. This process repeats until the end token is generated.
Pharmacophores
are abstract representations of the molecular features necessary for the molecular recognition of a ligand by a biological macromolecule, which are crucial for its biological activity. We selected the following pharmacophores types for one-hot encoding available in the RDKit (Landrum,, 2006) package: Hydrogen Bond Donors (HBD), Hydrogen Bond Acceptors (HBA), Aromatic Rings (AR), Hydrophobic Centers (HC), Positive Ionizable features (PIF), and Negative Ionizable features (NIF). In our approach, we represent pharmacophores through a one-hot encoding scheme. Specifically, we denote a pharmacophore feature vector as $\mathbf{p}=[p_{1},p_{2},\ldots,p_{n}]$, where each $p_{i}\in\{0,1\}$ of the pharmacophore type.
Data Preparation
To generate the data, we begin by selecting a set of building blocks and perform a sequence of $n$ random reactions from the available set, saving the resulting products. Next, we repeat the process but include the option to sample both the initial molecules and the previously generated products, effectively exploring the chemical space through random sampling. Once the molecules are sampled, we compute a random conformer for each one using Merck Molecular Force Field (MMFF) optimization from RDKit (Landrum,, 2006). Afterwards, we extract the pharmacophores and 3D coordinates of each molecule. This results in an input dataset where the pharmacophores and 3D coordinates serve as the inputs, and the output prediction corresponds to the synthetic tree.
3.2 Decoder Output: Molecules as Synthetic Trees
A molecule is generated by starting from the building blocks and iteratively applying the reactions. There are some cases where a reaction has multiple possible main products. For ease of discussion, we assume that reactions produce only one main product in this section.
As illustrated by the right panel of the first row in Figure 1, each node in the tree represents a compound, while the edges denote the reactions that transform one compound into another. This structure is useful for predicting feasible synthetic routes, optimizing reaction pathways, and planning the step-by-step construction of complex molecules.
For example, we begin with building block $B_{1}$, then we obtain the next building block, $B_{2}$ with the first reaction, $R_{1}$, that transforms $B_{1}$ into $P_{1}$. This reaction generates product $P_{1}$, thereafter we acquire the next building block, $B_{3}$, and the second reaction, $R_{2}$, which transforms $P_{1}$ into $P_{3}$. This process continues, step by step until the target molecule is reached, with each product serving as input for the next reaction, forming a complete synthetic tree.
Moreover, when sampling a batch, we can select any molecule as the final molecule and trace all the previous steps leading to the desired molecule, allowing for efficient exploration.
3.3 Architecture
SynthFormer is a two-part model that combines pharmacophore features with 3D spatial information from conformers to generate new molecules from building blocks. By integrating one-hot encoded pharmacophore vectors with conformer-derived coordinates, the EGNN encoder processes this data to generate rich embeddings. These embeddings are then passed to a transformer decoder that autoregressively generates building blocks and reactions, ensuring the synthesis of molecules with the desired pharmacophore features and synthetic accessibility.